# UNER: A Unified Prediction Head for Named Entity Recognition in Visually-rich Documents

### v1

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiadoi: 10.1145/3664647.3681473isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Information extractionccs: Applied computing Document analysisccs: Computing methodologies Transfer learning
@@ 4.1. Datasets
Fine-tuning: To demonstrate the effectiveness of our method, we perform fine-tuning on 7 popular VrD-NER benchmarks. Notably, when comparing with other prediction heads, we use the re-annotated versions of FUNSD (Jaume et al., 2019) and CORD (Park et al., 2019), namely FUNSD-r and CORD-r datasets (Zhang et al., 2023). The original datasets are not suitable for evaluating the ability to extract discontinuous entities, as the annotations in the original datasets are manually adjusted and do not reflect real-world scenarios. We use the re-annotated versions to better assess the entity extraction performance of different prediction heads. To provide solid experiments, we use another five benchmarks when comparing UNER-based models with other baselines. These benchmarks, including a cross-linguistic dataset in 7 languages, cover a wide range of document types, entity types, and languages.
Supervised Pre-training: During supervised pre-training, we expect the dataset to have a better diversity, so we choose two large datasets, SVRD and DocILE (Šimsa et al., 2023) for pre-training. The SVRD dataset originates from the ICDAR 2023 Competition [^1] https://rrc.cvc.uab.es/?ch=21&com=introduction and consists of densely annotated key-value linkings in various types of documents in real-world scenarios. By using the key entities as a description of the entity type, we transform these annotations into entity extraction labels, resulting in a VrD-NER dataset with hundreds of entity types. Since SVRD documents are primarily in Chinese, we utilize the DocILE dataset to enhance language diversity. This dataset comprises 5k documents with 55 entity types in English.
@@ 4.3. Effectiveness of UNER
Comparison with VrD-NER Baselines: The above results have proved the advantage of UNER compared to other prediction heads. In this section, in order to show that the UNER head can help in building excellent VrD-NER models, we further extended our comparison of two UNER-based models against other state-of-the-art VrD-NER models on 5 datasets: SROIE, EPHOIE, WildReceipt, SIBR, and XFUND.
The results of these comparisons are presented in Table-3. Note that XFUND is a cross-linguistic dataset and comprises seven sub-datasets, so we report the averaged F1 scores across them. The UNER head demonstrates highly competitive results on all datasets with both backbones, and it achieves the highest F1 score on three out of the five datasets overall, which is a significant achievement considering the number and diversity of the compared benchmarks. Notably, “LayoutLMv3+UNER” outperforms other methods by a significant margin (+$2.93\%$) on the WildReceipt dataset. These results further validate the effectiveness and compatibility of our UNER head to work with document transformer backbones.
@@ 4.4. Effectiveness of Supervised Pre-training
Comparison in Cross-linguistic Setting: Table-4 presents the detailed F1 scores of the sub-datasets in XFUND, a multi-linguistic dataset comprising 7 languages. The disparity in results between “LayoutMask+UNER” and “LayoutMask+UNER+SP” configurations highlight the efficacy of the supervised pre-training stage in enhancing the UNER-based model across all languages, which results in an average increase of $3.03\%$ in F1 score. Moreover, when compared to other methods, “LayoutMask+UNER+SP” achieves the highest performance on four sub-datasets: ZH, JA, FR, and DE. It is worth mentioning that the backbone model is pre-trained on English documents while the supervised pre-training datasets only include English and Chinese documents, so six out of the seven languages in XFUND are unseen before fine-tuning, which proves the effectiveness of supervised pre-training for cross-linguistic VrD-NER tasks.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681473isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Information extractionccs: Applied computing Document analysisccs: Computing methodologies Transfer learning
@@ 4.1. Datasets
Fine-tuning: To demonstrate the effectiveness of our method, we perform fine-tuning on 7 popular VrD-NER benchmarks. Notably, when comparing with other prediction heads, we use the re-annotated versions of FUNSD (Jaume et al., 2019) and CORD (Park et al., 2019), namely FUNSD-r and CORD-r datasets (Zhang et al., 2023). The original datasets are not suitable for evaluating the ability to extract discontinuous entities, as the annotations in the original datasets are manually adjusted and do not reflect real-world scenarios. We use the re-annotated versions to better assess the entity extraction performance of different prediction heads. To provide solid experiments, we use another five benchmarks when comparing UNER-based models with other baselines, which include a cross-linguistic dataset in 7 languages and cover a wide range of document types, entity types, and languages.
Supervised Pre-training: During supervised pre-training, we expect the dataset to have a better diversity, so we choose two large datasets, SVRD and DocILE (Šimsa et al., 2023) for pre-training. The SVRD dataset originates from the ICDAR 2023 Competition [^1] https://rrc.cvc.uab.es/?ch=21&com=introduction and consists of densely annotated key-value linkings in various types of documents. By using the key entities as a description of the entity type, we transform these annotations into entity extraction labels, resulting in a VrD-NER dataset with hundreds of entity types. Since SVRD documents are primarily in Chinese, we utilize the DocILE dataset to enhance language diversity, which comprises 5k documents with 55 entity types in English.
@@ 4.3. Effectiveness of UNER
Comparison with VrD-NER Baselines: The above results have proved the advantage of UNER compared to other prediction heads. In this section, in order to show that the UNER head can help in building excellent VrD-NER models, we further extended our comparison of two UNER-based models against other state-of-the-art VrD-NER models on 5 datasets: SROIE, EPHOIE, WildReceipt, SIBR, and XFUND. The results of these comparisons are presented in Table-3. Note that XFUND is a cross-linguistic dataset and comprises seven sub-datasets, so we report the averaged F1 scores across them. The UNER head demonstrates highly competitive results on all datasets with both backbones, and it achieves the highest F1 score on three out of the five datasets overall, which is a significant achievement considering the number and diversity of the compared benchmarks. Notably, “LayoutLMv3+UNER” outperforms other methods by a significant margin (+$2.93\%$) on the WildReceipt dataset. These results further validate the effectiveness and compatibility of our UNER head to work with document transformer backbones.
@@ 4.4. Effectiveness of Supervised Pre-training
Comparison in Cross-linguistic Setting: Table-4 presents the detailed F1 scores of the sub-datasets in XFUND, a multi-linguistic dataset comprising 7 languages. The disparity in results between “LayoutMask+UNER” and “LayoutMask+UNER+SP” configurations highlight the efficacy of the supervised pre-training stage in enhancing the UNER-based model across all languages, which results in an average increase of $3.03\%$ in F1 score. Moreover, when compared to other methods, “LayoutMask+UNER+SP” achieves the highest performance on four sub-datasets: ZH, JA, FR, and DE. It is worth mentioning that the backbone model is pre-trained on English documents while the supervised pre-training datasets only include English and Chinese documents, so six out of the seven languages in XFUND are unseen before fine-tuning, which proves the effectiveness of supervised pre-training for cross-linguistic tasks.